OpenStudy (anonymous):

I roll two fair six-sided dice. What is the probability that my roll totals 8?

OpenStudy (anonymous):

6/36

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Ok..... We have total of 36 cases when we roll 2 dice And we want the cases i which sum would be 8 So we have cases as : 2+6,3+5,4+4,4+4,5+3,6+2 = 6 cases (remember 4+4 cases would be 2 as the dices are not taken identical here for ex take a blue and other red so we can easily see there would be two seprate cases of 4+4) This probability=6/36=1/6

OpenStudy (anonymous):

thanks thats what i was thinking

OpenStudy (anonymous):

No problem! :-)